Express the rational expression in lowest terms
(3x+3)/(15x^2+24x=9)


Answer by benazir.sj@gmail.com(109)

(3x+3)/(15x^2+24x+9)
solution:
(3x+3) =3(x+1)
15x^2 +24x+9 =3(5x^2+8x+3)
=3(x+1)(5x+3) by factoring
(3x+3)/(15x^2+24x+9)

=3(x+1)/3(x+1)(5x+3)
=1/(5x+3)